# Chip-Scale Point-Source Sagnac Interferometer by Phase-Space Squeezing

### v1

@@
Matter-wave interferometry is essential to both science and technology. Phase-space squeezing has been shown to be an advantageous source of atoms, whereby the spread in momentum is decreased. Here, we show that the opposite squeezing may be just as advantageous. As a case in point, we analyze the effect of such a source on point source atom interferometry (PSI), which enables rotation sensing. We describe how a squeezed PSI (SPSI) increases the sensitivity and dynamic range while facilitating short cycle times and high repetition rates. We present regions in parameter space for which the figures of merit are improved by orders of magnitude and show that under some definition of compactness, the SPSI is superior by more than four orders of magnitude. The SPSI thus enables either enhancing the performance for standard size devices or maintaining the performance while miniaturizing to a chip-scale device, opening the door to real-life applications.
@@ Introduction
Here we propose a squeezed point-source interferometer (SPSI) for increasing the dynamic range of the interferometer and its sensitivity without increasing the time of operation or, alternatively, achieving the standard sensitivity with a smaller device. Moreso, a longstanding thought-after goal is to miniaturize rotation sensing  (?, ?, ?, ?), and we show that the SPSI opens the door to such miniaturization, and even enables a chip-scale device.
@@
Our method is based on adding a stage of pre-acceleration where the initial cloud goes through phase-space squeezing. Phase-space squeezing has been discussed extensively regarding delta-kick cooling  (?, ?, ?, ?, ?), where the spread in momentum is decreased at the expense of increasing the spread in position, while in contrast, here, we propose to increase the spread in momentum. As we show, integrating an inhomogeneous repulsive force to accelerate atom motion, either prior to or during the interferometer sequence, has the capability to substantially enhance the operational efficiency and sensitivity of a PSI device. We show that the figures of merit are enhanced by orders of magnitude, and under a definition of compactness, the enhancement in performance is about four orders of magnitude.
@@
An example of the effect of phase-space squeezing on the interferometer is demonstrated in Fig. 1. The squeezing leads to both an increase in the number of oscillations within the cloud size and an increase in the contrast.
@@
Figure 1: Principle of the squeezing effect in the interferometer. The atomic phase-space distribution is shown for the standard PSI interferometer (blue) and the SPSI interferometer (orange), at two different times: (A) at the start of the interferometer sequence, $t_{\rm{acc}}$, occurring just after the acceleration stage due to the repulsive-potential pulse, and (B) at the imaging time, $t_{\text{final}}$, following the complete interferometer sequence. The phase-space distribution at $t_{\rm{final}}$ becomes significantly tilted due to expansion, and the velocity-dependent interferometer phase introduces an oscillatory phase-space density upon detecting a single state. As shown at the bottom of (B), the latter oscillation produces a spatial density oscillation, which, upon being projected onto the position axis, forms the observed signal (output) of the interferometer. In the SPSI, when the initial phase-space distribution is squeezed with increased velocity, the subsequent expansion results in a phase-space distribution having a high aspect ratio. Upon projection onto the position axis, the SPSI signal shows significant improvement with more oscillations and an improved contrast.
Background
An established method for constructing a Laser-Pulse Atom Interferometry (LPAI) system involves a splitting a recombining laser giving rise to stimulated transitions (e.g., Raman, Bragg, or single photons for clock transitions) to implement a $\pi/2-\pi-\pi/2$ pulse sequence. The resulting phase shift (?), for free-falling atoms, is typically described by
where $\mathbf{K}_{\text{eff}}$ is the wave vector of the momentum given to the atom during the laser pulse, $\mathbf{a}$ is the constant acceleration of the platform relative to the atoms (including gravity), $T_{R}$ is the time interval between the interferometer pulses, $\mathbf{\Omega}$ represents the rotation of the platform, and $\mathbf{v}$ denotes the atom’s velocity. The second term in this equation is identified as the rotation phase. To effectively implement LPAI gyroscopes based on the $\pi/2-\pi-\pi/2$ pulse sequence, it is important to differentiate phase shifts induced by rotations from those caused by accelerations of the experimental apparatus. This can be achieved by combining the signals from two interferometers with counter-propagating atoms (?, ?, ?, ?).
The rotation phase in Eq. 1, can be expressed as the Sagnac phase (?) $\Delta\Phi_{\Omega}=(2m/\hbar)\mathbf{\Omega}\cdot\mathbf{A}$. In this context, $m$ is the atomic mass, while $\mathbf{A}$ is the area enclosed by the interferometer arms. Notably, these two expressions for the phase shift are equivalent, as demonstrated through the relation $\mathbf{A}=\left(\mathbf{v}_{r}T_{R}\right)\times\left(\mathbf{v}T_{R}\right)$, where $\mathbf{v}_{r}=\hbar\mathbf{K}_{\text{eff}}/m$ is the atom’s recoil velocity. Thus, the rotation measurement in LPAI gyroscopes depends on two distinct velocities: the recoil velocity and the axial velocity. The recoil velocity is linked to the momentum imparted on the atoms by the splitting and recombining laser. In the conventional Raman method, this momentum is equivalent to $2\hbar\mathbf{K}=\hbar\mathbf{K_{\text{eff}}}$, where $\mathbf{K}$ is the wave-vector of the Raman laser, although advanced techniques have demonstrated the potential to achieve significantly higher values  (?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?). The axial velocity is the topic of this work.
@@ Results
The point-source interferometer (PSI) is a specific kind of light-pulse atom interferometer in an expanding cloud of cold atoms  (?, ?, ?), and it is used in the following as a case study to examine in detail how the performance-enhancing phase-space squeezing works. The PSI method employs a single cloud of cold atoms and measures the rotation of the platform by probing the spatial frequency of the atomic density of a given output atomic state formed in the cloud. The method exploits the correlation between position and velocity generated by the expansion of the cloud throughout the interferometer sequence. When the final size of the cloud significantly exceeds its initial size, each atom’s final position, $\mathbf{x}$, is essentially dictated by its initial thermal velocity, $\mathbf{v}$. This allows for the approximation $\mathbf{x}\approx\mathbf{v}T_{\text{ex}}$, where $T_{\text{ex}}$ represents the total expansion time. In this limit, the Sagnac phase becomes  (?)
@@
According to Eq. 2, the probability of an atom with initial velocity $\mathbf{v}$ to be at the end of the interferometric sequence in a given internal state $|s\rangle$ is $P_{s}({\bf v})=1/2\cdot[1+\cos(\mathbf{k}_{v}\cdot\mathbf{v}+\phi_{s})]$, where $\mathbf{k}_{v}=2(\mathbf{K}_{\rm eff}\times\mathbf{\Omega})T_{R}^{2}$ and $\phi_{s}$ is determined by the gravitational acceleration. In an ideal case, the atomic cloud starts to expand freely from a point source, and the position of each atom is linearly proportional to its velocity $\mathbf{x}=\mathbf{v}T_{\rm{ex}}$, where $T_{\rm{ex}}$ is the time of expansion. In this case, the atoms form a spatial density that oscillates with a wave-vector $\mathbf{k}_{x}$, according to Eq. 2, from which one extracts the angular velocity component, which is transverse to the plane defined by the vectors $K_{\rm{eff}}$ and $k_{x}$.
@@
where $\sigma_{f}=\sqrt{\sigma_{x0}^{2}+T_{\rm{ex}}^{2}\sigma_{v0}}$ is the final cloud width, $\alpha=1-\sigma_{x0}^{2}/\sigma_{f}^{2}$, and the contrast $C$ is given by
@@
The sensitivity of an interferometer to rotation, limited by projection noise, hinges on two factors: the number of atoms per operation ($N$) and the frequency of operations ($\nu$). This sensitivity ($\delta\Omega$) is given by the expression  (?):
@@
In contrast to PSI, in interferometric rotation sensors, where the signal is atom population or light intensity in two output ports, the measured quantity is the Sagnac phase, which typically exhibits a linear correlation with the rotation. The smallest measurable phase difference determines the smallest detectable change in rotation, whether from zero rotation or from any given rotation. Hence, sensitivity characterizes both the smallest detectable rotation and the uncertainty in detecting a given rate of rotation. On the otherhand, in the PSI, the minimum detectable rotation by the interferometer is determined by the condition that the cloud size exceeds the period of spatial oscillation, expressed as $k_{x}\sigma_{f}>1$. In addition, the rotation should not surpass the threshold where the contrast significantly diminishes due to the ratio between the fringe periodicity and the initial cloud size, specified as $k_{x}\sigma_{x0}>1$. This implies a constraint on the fringe periodicity within the range $1/\sigma_{f}<k_{x}<1/\sigma_{x0}$, leading to a limitation on the dynamic range of the interferometer, namely, $\Omega_{\rm min}<\Omega<\Omega_{\rm max}$, where from Eq. 2 (and neglecting $\alpha$ from Eq. 3) we find
@@
such that $\Omega_{\rm max}/\Omega_{\rm min}\sim\sigma_{f}/\sigma_{x0}$. These limits are approximations closely related to the method used for image analysis. Previous works discussed this issue within various methods, including the ellipse fitting procedure  (?), phase shear  (?), or phase map  (?). Finally, it should be noted that by combining Eqs. 5 and 6, the sensitivity in detecting a rotation rate, $\delta\Omega$, is associated with the minimally detectable rate, $\Omega_{\rm{min}}$, by the relationship $\delta\Omega\approx 1/\sqrt{N/2}$. This suggests that for PSI, the sensitivity ($\delta\Omega$) solely pertains to uncertainty and may be significantly smaller (better) than the smallest detectable rotation rate.
Figure 2: Contrast and dynamic range. (A) Contrast vs. angular velocity ($\Omega$). Using 87Rb atoms, the simulation parameters are initial cloud size $\sigma_{x0}=100\,\rm{\mu m}$, temperature $T=5\,\rm{\mu K}$, and time between pulses $T_{R}=5\,\rm ms$. For the repulsive potential, we take a beam power of $P=1\,$W with a cross-section of $400\crossproduct 400\,\mu$m2, blue-detuned by $\Delta=2\pi\cdot 10\,$GHz, and focused such that in the direction of acceleration it gives rise to an inverted harmonic potential and in the transverse direction a constant potential. The harmonic profile was optimized for achieving a large squeezing factor before the atoms exceed the region where the repulsive potential is harmonic. Following Eqs. 8 and 11, and choosing $t_{\text{acc}}=80.6\,\mu$s, we find a squeezing parameter of 22.81. The purple line represents the analytical solution of PSI contrast, while the dashed orange line corresponds to SPSI, according to Eq. 4. The data points are the result of a numerical simulation. While the contrast reduction due to the ratio between the fringe periodicity and the initial cloud size determines the upper limit of the detection range, there exists a lower limit when the fringe period becomes larger than the final cloud size. These limits are roughly given by $\Omega_{\rm min}$ and $\Omega_{\rm max}$ (Eq. 6), which are presented in the graph as vertical dashed lines. It is evident that using the SPSI improves the contrast significantly, thereby increasing the detection dynamic range. At high angular velocities, the contrast of the data points decreases faster than the analytical solution due to the short spacing between fringes relative to the detection pixel size considered only in the numerical calculation. (B) The analytical curve of the detection range as a function of $\eta$ according to Eq. 6. The dashed line presents the SPSI simulated in (A) with $\eta=22.81$.
@@
The harmonic profile is quite advantageous for achieving a significant squeezing factor before the atoms reach the region with a repulsive potential. For example, for the parameters introduced in Fig. 2, $\omega_{\text{Harmonic}}=2\pi\cdot 1220\,$Hz.
The optical repulsive potential has the consequent effect of heating the atoms due to scattered photons  (?). Along the direction of beam propagation, the atoms gain momentum in correlation to the spatial intensity $I(\textbf{r})$, which can be compensated by shifting the initial cloud or, if necessary, by employing a counter-propagating beam. In addition, the scattering induces a random walk in velocity space in all directions, potentially enlarging the effective initial cloud size, $\tilde{\sigma}_{x0}$, and consequently decreasing the upper detection limit $\Omega_{\text{max}}$ (Eq. 6). Under the limitations we have taken for the laser power and typical parameter values, this effect reduces $\Omega_{\text{max}}$, e.g., by $\sqrt{2}$ for the parameters of Fig. 2. This effect can be reduced to a negligible level by improving the beam parameters, for example, by equally increasing the beam power $P$ and the detuning $\Delta$. More details about the heating process can be found in the appendix. One could also explore using different types of potential, such as magnetic gradients, to mitigate heating.
Figure 3: Relative sensitivity and measured angular velocity deviation. (A) Relative sensitivity $\delta\Omega/\Omega$ vs. $\Omega$, where $\Omega$ is the nominal (actual) angular velocity, for the same simulation parameters and notation as in Fig. 2. The analytical curves are determined according to Eq. 5. The data points for PSI are not shown at large angular velocities since $\delta\Omega/\Omega$ exceeds the value of 1, thus irrelevant. (B) Simulation results of the measured angular velocity deviation $\Delta\Omega\equiv(\Omega_{\rm meas}-\Omega)$ for different angular velocities. Here $\Delta\Omega$ denotes the discrepancy between the measured and nominal angular velocities, while the sensitivity (uncertainty) $\delta\Omega$ is represented by the error bars. The data point labels are the same as in (A), while the black dashed line illustrates the point-source (PS) limit of zero deviation. The data points for PSI are not shown at large angular velocities since $\Delta\Omega$ is too large. The enhanced performance of the SPSI in dynamic range, precision, and sensitivity (base noise) is clearly visible.
@@
To estimate the performance of the SPSI and compare it to that of the standard PSI, we calculated its characteristics analytically and numerically (see Methods for the latter).
@@
For the comparison, under the same initial conditions (temperature, cloud size, and number of atoms), the ratio between the sensitivities obtained in the two methods is $\delta\Omega_{SPSI}/\delta\Omega_{PSI}\approx(C_{PSI}/C_{SPSI})/\eta$. Within the dynamic range of both methods, the contrast ratio is approximately $\sim 1$, while the detection limits ratio is $\Omega_{\rm min,SPSI}/\Omega_{\rm min,PSI}=1/\eta$ and $\Omega_{\rm max,SPSI}/\Omega_{\rm max,PSI}=\eta$. Consequently, for chip-scale devices, we can expect a sensitivity improvement of about $\eta$ and an increase in the dynamic range $\Omega_{\rm max}/\Omega_{\rm min}$ by a factor of $\eta^{2}$.
Specifically, we consider the following chip-scale scenario: a rectangular vacuum cell with dimensions of $1\,\rm{mm}\times 10\,\rm{mm}\times 10\,\rm{mm}$, with the short dimension aligned perpendicular to the direction of the repulsive potential’s acceleration. The SPSI sequence commences with the capture of an atomic cloud comprising $N=10^{6}$ 87Rb atoms, with a radius of $\sigma_{x0}=100\,\rm{\mu m}$ and temperature of $T=2\,\rm{\mu K}$, using a magneto-optical trap and optical molasses setup. Employing a fully harmonic potential characterized by dimensions of $x_{0}=200\,\rm{\mu m}$ and $z_{0}=400\,\rm{\mu m}$, with total power of $P=1\,\rm{W}$ and an acceleration time of $t_{\rm acc}=0.1\,\rm{ms}$, yields a squeezing parameter $\eta=46.2$, potential frequency $\omega_{\text{Harmonic}}=2\pi\cdot 1220\,$Hz, and final velocity uncertainty $\tilde{\sigma}_{v0}=0.64\,$m/s. Following acceleration, the atoms undergo a Ramsey sequence with a time $T_{R}=2.5\,\rm{ms}$, resulting in a total fall height of $h=120\,\rm{\mu m}$. The final cloud dimensions are $\sigma_{f,z}=170\,\rm{\mu m}$ height and $\sigma_{f,x}=0.33\,\rm{cm}$ width. Without heating, a vacuum dimension of $0.5\,\rm{mm}$ would suffice, but if heating is not mitigated, a vacuum dimension of $1\,\rm{mm}$ would be required, or alternatively, one would have to take into account that a significant amount of atoms could not be recaptured for the next cycle, and reloading atoms from a source would be required.

### v2

@@
Matter-wave interferometry is essential to both science and technology. Phase-space squeezing has been shown to be an advantageous source of atoms, whereby the spread in momentum is decreased. Here, we show that the opposite squeezing may be just as advantageous. As an exemplification, we analyze the effect of such a source on point source atom interferometry (PSI), which enables rotation sensing. We describe how a squeezed PSI (SPSI) increases the sensitivity and dynamic range while facilitating short cycle times and high repetition rates. We present regions in parameter space for which the figures of merit are improved by orders of magnitude and show that under some definition of compactness, the SPSI is superior by more than four orders of magnitude. The SPSI thus enables either enhancing the performance for standard size devices or maintaining the performance while miniaturizing to a chip-scale device, opening the door to real-life applications.
@@ Introduction
Here we propose a squeezed point-source interferometer (SPSI) for increasing the dynamic range of the interferometer and its sensitivity without increasing the time of operation or, alternatively, achieving the standard sensitivity with a smaller device. Moreso, a longstanding sought-after goal is to miniaturize rotation sensing  (?, ?, ?, ?), and we show that the SPSI opens the door to such miniaturization, and even enables a chip-scale device.
@@
Our method is based on adding a stage of pre-acceleration where the initial cloud goes through phase-space squeezing. Phase-space squeezing has been discussed extensively regarding delta-kick cooling  (?, ?, ?, ?, ?), in which the spread in momentum is decreased at the expense of increasing the spread in position. In contrast, here, we propose to increase the spread in momentum while the spread in position is effectively reduced. As we show, integrating an inhomogeneous repulsive force to accelerate atom motion, either prior to or during the interferometer sequence, has the capability to substantially enhance the operational efficiency and sensitivity of a PSI device. We show that the figures of merit are enhanced by orders of magnitude, and under a definition of compactness, the enhancement in performance is about four orders of magnitude.
@@
An example of the effect of phase-space squeezing on the interferometer is illustrated in Fig. 1. The squeezing leads to both an increase in the number of oscillations within the cloud size and an increase in the contrast.
@@
Figure 1: Principle of the squeezing effect in the interferometer. The atomic phase-space distribution is shown for the standard PSI interferometer (blue) and the SPSI interferometer (orange), at two different times: (A) at the start of the interferometer sequence, $t_{\rm{acc}}$, occurring just after the acceleration stage due to the repulsive-potential pulse, and (B) at the imaging time, $t_{\text{final}}$, following the complete interferometer sequence. The phase-space distribution at $t_{\rm{final}}$ becomes tilted due to expansion, and the velocity-dependent interferometer phase introduces an oscillatory phase-space density upon detecting a single state. As shown at the bottom of (B), the latter oscillation produces a spatial density oscillation, which, upon being projected onto the position axis, forms the observed signal (output) of the interferometer. In the SPSI, when the initial phase-space distribution is squeezed with increased velocity, the subsequent expansion results in a phase-space distribution having a high aspect ratio. Upon projection onto the position axis, the SPSI signal shows improvement with more oscillations and an improved contrast.
@@ Results
An established method for constructing a Laser-Pulse Atom Interferometry (LPAI) system involves a splitting and recombining laser giving rise to stimulated transitions (e.g., Raman, Bragg, or single photons for clock transitions) to implement a $\pi/2-\pi-\pi/2$ pulse sequence. The resulting phase shift (?), for free-falling atoms, is typically described by
where $\mathbf{K}_{\text{eff}}$ is the wave vector of the momentum given to the atom during the laser pulse, $\mathbf{a}$ is the constant acceleration of the platform relative to the atoms (including gravity), $T_{R}$ is the time interval between the interferometer pulses, $\mathbf{\Omega}$ represents the rotation rate of the platform, and $\mathbf{v}$ denotes the initial velocity of the atom. The second term in this equation is identified as the rotation phase. To effectively implement LPAI gyroscopes based on the $\pi/2-\pi-\pi/2$ pulse sequence, it is important to differentiate phase shifts induced by rotations from those caused by accelerations of the experimental apparatus. This can be achieved by combining the signals from two interferometers with counter-propagating atoms (?, ?, ?, ?).
The rotation phase in Eq. 1, can be expressed as the Sagnac phase (?) $\Delta\Phi_{\Omega}=(2m/\hbar)\mathbf{\Omega}\cdot\mathbf{A}$. In this context, $m$ is the atomic mass, while $\mathbf{A}$ is the area enclosed by the interferometer arms. Notably, these two expressions for the phase shift are equivalent, as demonstrated through the relation $\mathbf{A}=\left(\mathbf{v}_{r}T_{R}\right)\times\left(\mathbf{v}T_{R}\right)$, where $\mathbf{v}_{r}=\hbar\mathbf{K}_{\text{eff}}/m$ is the atom’s recoil velocity. Thus, the rotation measurement in LPAI gyroscopes depends on two distinct velocities: the recoil velocity and the initial velocity of the atom. The recoil velocity is linked to the momentum imparted on the atoms by the splitting and recombining laser. In the conventional Raman method, this momentum is equivalent to $2\hbar\mathbf{K}=\hbar\mathbf{K_{\text{eff}}}$, where $\mathbf{K}$ is the wave-vector of the Raman laser, although advanced techniques have demonstrated the potential to achieve notable higher values  (?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?). The position and velocity distribution of the atoms prior to and during the interferometer sequence is the topic of this work.
The point-source interferometer (PSI) is a specific kind of light-pulse atom interferometer in an expanding cloud of cold atoms  (?, ?, ?), and it is used in the following as a case study to examine in detail how the performance-enhancing phase-space squeezing works. The PSI method employs a single cloud of cold atoms and measures the rotation of the platform by probing the spatial frequency of the atomic density of a given output internal atomic state. The method exploits the correlation between position and velocity generated by the expansion of the cloud throughout the interferometer sequence. When the final cloud becomes much larger than the initial cloud, each atom’s final position, $\mathbf{x}$, is essentially dictated by its initial thermal velocity, $\mathbf{v}$. This allows for the approximation $\mathbf{x}\approx\mathbf{v}T_{\text{ex}}$, where $T_{\text{ex}}$ represents the total expansion time. In this limit, the Sagnac phase becomes  (?)
@@
According to Eq. 2, the probability of an atom with initial velocity $\mathbf{v}$ to be at the end of the interferometric sequence in a given internal state $|s\rangle$ is $P_{s}({\bf v})=1/2\cdot[1+\cos(\mathbf{k}_{v}\cdot\mathbf{v}+\phi_{s})]$, where $\mathbf{k}_{v}=2(\mathbf{K}_{\rm eff}\times\mathbf{\Omega})T_{R}^{2}$ and $\phi_{s}$ is determined by gravity and platform acceleration. In an ideal case, the atomic cloud starts to expand freely from a point source, and the position of each atom is linearly proportional to its velocity $\mathbf{x}=\mathbf{v}T_{\rm{ex}}$, where $T_{\rm{ex}}$ is the time of expansion. In this case, the atoms form a spatial density that oscillates with a wave-vector $\mathbf{k}_{x}$, according to Eq. 2, from which one extracts the angular velocity component, which is transverse to the plane defined by the vectors $K_{\rm{eff}}$ and $k_{x}$.
@@
where $\sigma_{f}=\sqrt{\sigma_{x0}^{2}+T_{\rm{ex}}^{2}\sigma_{v0}^{2}}$ is the final cloud width, $\alpha=1-\sigma_{x0}^{2}/\sigma_{f}^{2}$, and the contrast $C$ is given by
@@
The sensitivity of an interferometer to rotation, limited by projection noise, hinges on two additional factors: the number of atoms per operation ($N$) and the frequency of operations ($\nu$). This sensitivity ($\delta\Omega$) is given by the expression  (?):
@@
In contrast to PSI, in interferometric rotation sensors, where the signal is atom population or light intensity in two output ports, the measured quantity is the Sagnac phase, which typically exhibits a linear correlation with the rotation. The smallest measurable phase difference determines the smallest detectable change in rotation, whether from zero rotation or from any given rotation. Hence, sensitivity characterizes both the smallest detectable rotation and the uncertainty in detecting a given rate of rotation. On the other hand, in the PSI, the minimum detectable rotation by the interferometer is determined by the condition that the cloud size exceeds the period of spatial oscillation, expressed as $k_{x}\sigma_{f}>1$. In addition, the contrast diminishes if the rotation rate exceeds the threshold where the fringe period becomes larger than the initial cloud size, namely $k_{x}\sigma_{x0}>1$. This implies a constraint on the fringe periodicity within the range $1/\sigma_{f}<k_{x}<1/\sigma_{x0}$, leading to a limitation on the dynamic range of the interferometer, namely, $\Omega_{\rm min}<\Omega<\Omega_{\rm max}$, where from Eq. 2 (and neglecting $\alpha$ from Eq. 3) we find
@@
such that $\Omega_{\rm max}/\Omega_{\rm min}\sim\sigma_{f}/\sigma_{x0}$. These limits are approximations closely related to the method used for image analysis. Previous works discussed this issue within various methods, including the ellipse fitting procedure  (?), phase shear  (?), or phase map  (?). Finally, it should be noted that by combining Eqs. 5 and 6, the sensitivity in detecting a rotation rate, $\delta\Omega$, is associated with the minimally detectable rate, $\Omega_{\rm{min}}$, by the relationship $\delta\Omega\approx 1/\sqrt{N/2}$. This suggests that for PSI, the sensitivity ($\delta\Omega$) solely pertains to uncertainty and may be much smaller (better) than the smallest detectable rotation rate.
Figure 2: Contrast and dynamic range. (A) Contrast vs. angular velocity ($\Omega$). Using 87Rb atoms, the simulation parameters are initial cloud size $\sigma_{x0}=100\,\rm{\mu m}$, temperature $T=5\,\rm{\mu K}$, and time between pulses $T_{R}=5\,\rm ms$. For the repulsive potential, we take a beam power of $P=1\,$W with a cross-section of $400\crossproduct 400\,\mu$m2, blue-detuned by $\Delta=2\pi\cdot 10\,$GHz, and focused such that in the direction of acceleration it gives rise to an inverted harmonic potential and in the transverse direction a constant potential. The harmonic profile was optimized for achieving a large squeezing factor before the atoms exceed the region where the repulsive potential is harmonic. Following Eqs. 8 and 11, and choosing $t_{\text{acc}}=80.6\,\mu$s, we find a squeezing parameter of 22.81. The purple line represents the analytical solution of PSI contrast, while the dashed orange line corresponds to SPSI, according to Eq. 4. The data points are the result of a numerical simulation. While the contrast reduction due to the ratio between the fringe periodicity and the initial cloud size determines the upper limit of the detection range, there exists a lower limit when the fringe period becomes larger than the final cloud size. These limits are roughly given by $\Omega_{\rm min}$ and $\Omega_{\rm max}$ (Eq. 6), which are presented in the graph as vertical dashed lines. It is evident that using the SPSI improves the contrast greatly, thereby increasing the detection dynamic range. At high angular velocities, the contrast of the data points decreases faster than the analytical solution due to the short spacing between fringes relative to the detection pixel size considered only in the numerical calculation. (B) The analytical curve of the detection range as a function of $\eta$ according to Eq. 6. The dashed line presents the SPSI simulated in (A) with $\eta=22.81$.
@@
The harmonic profile is quite advantageous for achieving a large squeezing factor before the atoms reach the region with a repulsive potential. For example, for the parameters introduced in Fig. 2, $\omega_{\text{Harmonic}}=2\pi\cdot 1220\,$Hz.
The optical repulsive potential has the consequent effect of heating the atoms due to scattered photons  (?). Along the direction of beam propagation, the atoms gain momentum in correlation to the spatial intensity $I(\textbf{r})$, which can be compensated by shifting the initial cloud position or, if necessary, by employing a counter-propagating beam. In addition, the scattering induces a random walk in velocity space in all directions, potentially enlarging the effective initial cloud size, $\tilde{\sigma}_{x0}$, and consequently decreasing the upper detection limit $\Omega_{\text{max}}$ (Eq. 6). Under the limitations we have taken for the laser power and typical parameter values, this effect reduces $\Omega_{\text{max}}$, e.g., by $\sqrt{2}$ for the parameters of Fig. 2. This effect can be reduced to a negligible level by improving the beam parameters, for example, by equally increasing the beam power $P$ and the detuning $\Delta$. More details about the heating process can be found in the appendix. One could also explore using different types of potential, such as magnetic gradients, to mitigate heating.
Figure 3: Relative sensitivity and measured angular velocity deviation. (A) Relative sensitivity $\delta\Omega/\Omega$ vs. $\Omega$, where $\Omega$ is the nominal (actual) angular velocity, for the same simulation parameters and notation as in Fig. 2. The analytical curves are determined according to Eq. 5. The data points for PSI are not shown at large angular velocities since $\delta\Omega/\Omega$ exceeds the value of 1, thus irrelevant. (B) Simulation results of the measured angular velocity deviation $\Delta\Omega\equiv(\Omega_{\rm meas}-\Omega)$ for different angular velocities. Here $\Delta\Omega$ denotes the discrepancy between the measured and nominal angular velocities, while the sensitivity (uncertainty) $\delta\Omega$ is represented by the error bars. The data point labels are the same as in (A), while the black dashed line illustrates the point-source (PS) limit of zero deviation. The data points for PSI are not shown at large angular velocities since $\Delta\Omega$ is too large. The enhanced performance of the SPSI in dynamic range and sensitivity (measurement uncertainty) is clearly visible.
@@
To estimate the performance of the SPSI and compare it to that of the standard PSI, we calculated its characteristics analytically and numerically (see Materials and Methods for the latter).
@@
For the comparison, under the same initial conditions (temperature, cloud size, and number of atoms), the ratio between the sensitivities obtained in the two methods is $\delta\Omega_{SPSI}/\delta\Omega_{PSI}\approx(C_{PSI}/C_{SPSI})/\eta$. Within the dynamic range of both methods, the contrast ratio is approximately $\sim 1$, while the detection limits ratio is $\Omega_{\rm min,SPSI}/\Omega_{\rm min,PSI}=1/\eta$ and $\Omega_{\rm max,SPSI}/\Omega_{\rm max,PSI}=\eta$. Consequently, for chip-scale devices, we can expect the sensitivity improvement of about $\eta$ and an increase in the dynamic range $\Omega_{\rm max}/\Omega_{\rm min}$ by a factor of $\eta^{2}$.
Specifically, we consider the following chip-scale scenario: a rectangular vacuum cell with dimensions of $1\,\rm{mm}\times 10\,\rm{mm}\times 10\,\rm{mm}$, with the short dimension aligned perpendicular to the direction of the repulsive potential’s acceleration. The SPSI sequence commences with the capture of an atomic cloud comprising $N=10^{6}$ 87Rb atoms, with a radius of $\sigma_{x0}=100\,\rm{\mu m}$ and temperature of $T=2\,\rm{\mu K}$, using a magneto-optical trap and optical molasses setup. Employing a fully harmonic potential characterized by dimensions of $x_{0}=200\,\rm{\mu m}$ and $z_{0}=400\,\rm{\mu m}$, with total power of $P=1\,\rm{W}$ and an acceleration time of $t_{\rm acc}=0.1\,\rm{ms}$, yields a squeezing parameter $\eta=46.2$, potential frequency $\omega_{\text{Harmonic}}=2\pi\cdot 1220\,$Hz, and final velocity uncertainty $\tilde{\sigma}_{v0}=0.64\,$m/s. Following acceleration, the atoms undergo a Ramsey sequence with a time $T_{R}=2.5\,\rm{ms}$, resulting in a total fall height of $h=120\,\rm{\mu m}$. The final cloud dimensions are $\sigma_{f,z}=170\,\rm{\mu m}$ height and $\sigma_{f,x}=0.33\,\rm{cm}$ width. Without heating, a vacuum dimension of $0.5\,\rm{mm}$ would suffice, but if heating is not mitigated, a vacuum dimension of $1\,\rm{mm}$ would be required, or alternatively, one would have to take into account that a considerable amount of atoms could not be recaptured for the next cycle, and reloading atoms from a source would be required.